# Modeling Non-Gaussianities in Pulsar Timing Array data analysis using Gaussian Mixture Models

### v1

@@ II.1 Gaussian noise
$\ln p(\vec{a},\vec{\lambda}|\delta t)=$ $-\frac{1}{2}\left[\delta t-\bm{F}a\right]^{\top}\bm{N}^{-1}\left[\delta t-\bm{F}a\right]-\frac{1}{2}\ln\det 2\pi\bm{N}$ (2)
$-\frac{1}{2}a^{\top}\bm{\Phi}^{-1}a-\frac{1}{2}\ln\det 2\pi\bm{\Phi}$
$+\ln\Pi(\vec{\lambda}),$
@@
$\ln\int p(\vec{a},\vec{\lambda}|\delta t)d\vec{a}$ $=-\frac{1}{2}\delta t^{\top}\bm{C}^{-1}\delta t-\frac{1}{2}\ln\det\{2\pi\bm{C}\}+\ln\Pi(\vec{\lambda})$ (3)
$=\ln\mathcal{L}(\delta t|\vec{\lambda})+\ln\Pi(\vec{\lambda}),$
@@ II.2 Non-Gaussian noise
$p_{a}(a|\Phi,\vec{\alpha},\vec{\mu},\vec{c})$ $=\sum_{i=1}^{N}\alpha_{i}\frac{\exp\left\{-\frac{1}{2}\left[a-\mu_{i}\right]^{2}(c_{i}\Phi)^{-1}\right\}}{\sqrt{2\pi(c_{i}\Phi)}}$ (4)
$=\sum_{i=1}^{N}\alpha_{i}\varphi(a,\mu_{i},\sqrt{c_{i}\Phi}),$
@@
$\ln\int p(a,\Phi,\vec{\alpha},\vec{c}|\delta t)da=$ $\ln\sum_{i=1}^{N}\alpha_{i}\mathcal{L}(\delta t|c_{i}\Phi)$ (5)
$+\ln\Pi(\Phi)+\ln\Pi(\vec{c})+\ln\Pi(\vec{\alpha}),$
@@ II.4 Non-Gaussian Likelihood
$p_{n}(a|\Phi_{n},\alpha,c)$ $=(1-\alpha)\times\varphi(a,0,\sqrt{\Phi_{n}})+\alpha\times\varphi(a,0,\sqrt{c\Phi_{n}}),$ (13)
@@
$\mathcal{L}(\delta t|\Phi^{c}_{n},\Phi^{s}_{n},\alpha,c,\vec{\lambda})$ $=(1-\alpha)^{2}\mathcal{L}(\delta t|\Phi_{n},\Phi_{n},\vec{\lambda})$ (14)
$+\alpha^{2}\mathcal{L}(\delta t|c\Phi_{n},c\Phi_{n},\vec{\lambda})$
$+(1-\alpha)\alpha\mathcal{L}(\delta t|\Phi_{n},c\Phi_{n},\vec{\lambda})$
$+\alpha(1-\alpha)\mathcal{L}(\delta t|c\Phi_{n},\Phi_{n},\vec{\lambda}).$
@@ II.5 Free spectrum
When $a_{n}$ is described as a GMM, we found no closed form expression of the ensemble average for the sum of the squares. However, the result can be easily obtained numerically by sampling the GMM. For a set of non-Gaussianity parameters $\Phi_{n}$, $\alpha$ and $c$, given that we have inferred the posterior distribution $p(\Phi_{n},\alpha,c|\delta t)$ from data $\delta t$, the probability distribution $p(a_{n},\Phi_{n},\vec{\alpha},\vec{c}|\delta t)$ is given by
@@
$p(a_{n}|\delta t)$ $=\int p(a_{n},\Phi_{n},\alpha,c|\delta t)d\Phi_{n}d\alpha dc$ (17)
$\approx\frac{1}{N_{D}}\sum_{j=1}^{N_{D}}p(a_{n}|\Phi_{j},\alpha_{j},c_{j}).$
@@
where the $a_{nj,i}^{c},a_{nj,i}^{s}$ are sampled from $p(a_{n},\Phi_{nj},\alpha_{j},c_{j}|\delta t)$ and $N_{s}$ is the total number of samples used to numerically estimate the average for the $j$-th posterior sample of $p(\Phi_{n},\alpha,c|\delta t)$. Here, we discard uneven moments because we only consider symmetric distributions.
@@ III Results
For each pulsar we consider a timing model accounting for the spin-down rate of that pulsar described by a quadratic polynomial in time and we marginalize over the first order errors in the timing model parameters [14, 19].
@@ III.1 Single pulsar noise
We test the model by injecting a time-correlated noise that has a consistent statistic across all frequency bins. In this example, we consider a noise that has Fourier coefficients $a$ distributed accroding to GMM of 2 Gaussians with $\alpha=0.5$ and $c=3$, following a powerlaw as in Equation 20 with $\log_{10}A=-13.5$ and $\gamma=3$ across 50 frequency bins that are the 50 first harmonics of $1/T_{obs}$. Recent works have shown the advantages of directly sampling the Fourier coefficients [24, 29]. The only disadvantange comes from the high number of coefficients there are to sample, but the recovered posterior for the unmarginalized parameters are identical to the traditionally marginalized likelihood.
The result of one simulation is shown in Figure 2 to verify that we indeed recover the injected values of the parameters. We only show the hyperparameters characterizing the noise, while numerically marginalizing over all Fourier coefficients by sampling them. We have used 50 frequencies, for a total of 100 Fourier coefficients that we fit for. This approach is straightforward as it requires no analytical marginalization. However, the sampling is affected by the high dimensionality of the parameter space and many samples are required to converge to the proper distribution.
@@ III.2 Common red noise
Figure 3: Bayes factor $\mathcal{B}^{NG}_{G}$ between the non-Gaussian and Gaussian models as a function of the injected relative excess kurtosis $\Delta\bar{M}_{4}$ calculated using Equation 7 for many simulations. The solid line is the median and the shaded region is bounded by the 16th and 84th percentiles.
@@ III.3 Background plus few sources
Figure 7: Excess moments as a function of the averaged SNR of the source for $N_{p}$ pulsars with $10$ years of observation, $100$ ns noise level and a GWB signal with $h_{c}=10^{-15}(f/f_{\rm yr})^{-2/3}$. Since $f_{\rm gw}=3/(10\textrm{yr})$, the source produces 3 cycles in the data. This plot shows the median excess moments estimated from a 100 realization of the CGW parameters. The shaded colored areas show the 16th and 84th percentiles of the realizations. Only the amplitude of the CGW was fixed to control the SNR.
@@ IV Conclusion
In this work, we introduced a method to model non-Gaussian features in PTA data by using GMMs as priors for the Fourier coefficients of stochastic processes. This approach enables analytical marginalization of the likelihood, making it computationally efficient and directly compatible with existing PTA analysis pipelines. We validated the method through simulations involving both single pulsar noise and common red noise processes, demonstrating that the model can robustly identify departures from Gaussianity using Bayesian model selection. The Bayes factor increases significantly in the presence of excess kurtosis, confirming the sensitivity of the approach to higher-order statistical structure. Furthermore, we applied the method to a more astrophysically motivated scenario involving a superposition of a Gaussian GWB and a few CGW sources. In this case, we showed that non-Gaussian features become detectable when a small number of loud sources dominate the signal, and that these features are suppressed once the deterministic CGWs are correctly modeled and subtracted. When a realistic GWB is generated from the superposition of the individual CGW emissions of a simulated population of SMBHBs, the non-Gaussianities are harder to detect and might require very large arrays of pulsars to increase their detectability.
@@ Acknowledgements.
We acknowledge support from the European Union’s H2020 ERC Advanced Grant “PINGU” (Grant No. 101142079) and from the PRIN 2022 grant ”GRAPE” (CUP H53D2300091 0006).
@@ Appendix A Moments of a Gaussian mixture
$M_{n}$ $=\int(x-\alpha\mu_{0})^{n}p(x)dx$ (25)
$=(1-\alpha)\int(x-\alpha\mu_{0})^{n}\varphi(x,0,\sqrt{\Phi})dx+\alpha\int(x-\alpha\mu_{0})^{n}\varphi(x,\mu_{0},\sqrt{c\Phi})dx$
$=(1-\alpha)\int\tilde{x}^{n}\varphi(\tilde{x},-\alpha\mu_{0},\sqrt{\Phi})d\tilde{x}+\alpha\int\tilde{x}^{n}\varphi(\tilde{x},(1-\alpha)\mu_{0},\sqrt{c\Phi})d\tilde{x}$
$=(1-\alpha)\mathcal{M}_{n}(-\alpha\mu_{0},\sqrt{\Phi})+\alpha\mathcal{M}_{n}((1-\alpha)\mu_{0},\sqrt{c\Phi}),$
@@ Appendix B Marginalized Gaussian likelihood
$-\frac{1}{2}a^{\top}\Sigma a+a^{\top}\left[F^{\top}N^{-1}\delta t+\Phi^{-1}\mu\right]$ (28)
$=$ $-\frac{1}{2}\left[a-\Sigma^{-1}F^{\top}N^{-1}\delta t-\Sigma^{-1}\Phi^{-1}\mu\right]^{\top}\Sigma\left[a-\Sigma^{-1}F^{\top}N^{-1}\delta t-\Sigma^{-1}\Phi^{-1}\mu\right]$
$+\frac{1}{2}\left[\mu^{\top}\Phi^{-1}\Sigma^{-1}+\delta t^{\top}N^{-1}F\Sigma^{-1}\right]\Sigma\left[\Sigma^{-1}F^{\top}N^{-1}\delta t+\Sigma^{-1}\Phi^{-1}\mu\right]$
@@
[…]
$\times\exp\left\{-\frac{1}{2}\mu^{\top}\left[\Phi^{-1}-\Phi^{-1}\Sigma^{-1}\Phi^{-1}\right]\mu+\mu^{\top}\Phi^{-1}\Sigma^{-1}F^{\top}N^{-1}\delta t\right\}$
@@
$\mathcal{L}(\delta t|\Phi,\mu)$ $=\int_{-\infty}^{+\infty}da\mathcal{L}(\delta t|a,\Phi,\mu)$ (30)
$=\frac{\exp\left\{-\frac{1}{2}\delta t^{\top}\left[N+F^{\top}\Phi F\right]^{-1}\delta t\right\}}{|2\pi N|^{1/2}|\Phi|^{1/2}|\Sigma|^{1/2}}$
$\times\exp\left\{-\frac{1}{2}\mu^{\top}\left[\Phi^{-1}-\Phi^{-1}\Sigma^{-1}\Phi^{-1}\right]\mu+\mu^{\top}\Phi^{-1}\Sigma^{-1}F^{\top}N^{-1}\delta t\right\}$
@@ Appendix C Gram-Charlier A expansion for kurtosis
$H_{IJKL}(\vec{a})=$ $(-1)^{4}\ \Pi(\vec{a})^{-1}\frac{\partial^{4}}{\partial a_{I}\partial a_{J}\partial a_{K}\partial a_{L}}\Pi(\vec{a})$ (34)
$=$ $\Phi^{-4}\left[(\sum_{i}\chi_{iI}^{-1}a_{i})(\sum_{j}\chi_{jJ}^{-1}a_{j})(\sum_{k}\chi_{kK}^{-1}a_{k})(\sum_{l}\chi_{lL}^{-1}a_{l})\right]$
$-\Phi^{-3}\left[\chi^{-1}_{IJ}(\sum_{k}\chi_{kK}^{-1}a_{k})(\sum_{l}\chi_{lL}^{-1}a_{l})+\chi^{-1}_{IK}(\sum_{j}\chi_{jJ}^{-1}a_{j})(\sum_{l}\chi_{lL}^{-1}a_{l})+\chi^{-1}_{IL}(\sum_{j}\chi_{jJ}^{-1}a_{j})(\sum_{k}\chi_{kK}^{-1}a_{k})\right$.
$-\left.\chi^{-1}_{JK}(\sum_{i}\chi_{iI}^{-1}a_{i})(\sum_{l}\chi_{lL}^{-1}a_{l})+\chi^{-1}_{JL}(\sum_{i}\chi_{iI}^{-1}a_{i})(\sum_{k}\chi_{kK}^{-1}a_{k})+\chi^{-1}_{KL}(\sum_{i}\chi_{iI}^{-1}a_{i})(\sum_{j}\chi_{jJ}^{-1}a_{j})\right]$
$+\Phi^{-2}\bigg{[}\chi^{-1}_{IJ}\chi^{-1}_{KL}+\chi^{-1}_{IK}\chi^{-1}_{JL}+\chi^{-1}_{IL}\chi^{-1}_{JK}\bigg{]}.$
@@
$\mathcal{L}(\delta t|\Phi,\kappa)$ $=\int_{-\infty}^{+\infty}da\mathcal{L}(\delta t|a,\Phi,\kappa,\mu=0)$ (35)
$=\mathcal{C}^{-1}\times\mathcal{L}(\delta t|\Phi,\kappa=0)\int da\frac{\exp\left\{-\frac{1}{2}\left[a-\hat{a}\right]^{\top}\Sigma\left[a-\hat{a}\right]\right\}}{|2\pi\Sigma^{-1}|}\left[1+\frac{1}{4!}\sum_{IJKL}\kappa_{IJKL}H_{IJKL}(\vec{a}_{n})\right]$
$=\mathcal{C}^{-1}\times\mathcal{L}(\delta t|\Phi,\kappa=0)\left[1+\frac{1}{24}\sum_{IJKL}\kappa_{IJKL}\textrm{E}_{\Sigma}\left\{H_{IJKL}\right\}\right]$
@@
$\textrm{E}_{\Sigma}\left\{H_{IJKL}\right\}$ $=\Phi^{-4}\sum_{ijkl}\chi^{-1}_{iI}\chi^{-1}_{jJ}\chi^{-1}_{kK}\chi^{-1}_{lL}\textrm{E}_{\Sigma}\left\{a_{i}a_{j}a_{k}a_{l}\right\}$ (36)
$-\Phi^{-3}\left[\chi^{-1}_{IJ}\sum_{kl}\chi^{-1}_{kK}\chi^{-1}_{lL}\textrm{E}_{\Sigma}\left\{a_{k}a_{l}\right\}+\textrm{other combinations}\right]$
$+\Phi^{-2}\left[\chi^{-1}_{IJ}\chi^{-1}_{KL}+\textrm{other combinations}\right].$
@@
$\textrm{E}_{\Sigma}\left\{a_{i}a_{j}a_{k}a_{l}\right\}$ $=\hat{a}_{i}\hat{a}_{j}\hat{a}_{k}\hat{a}_{l}$ (37)
$+\hat{a}_{i}\hat{a}_{j}\Sigma_{kl}^{-1}+\hat{a}_{i}\hat{a}_{k}\Sigma_{jl}^{-1}+\hat{a}_{i}\hat{a}_{l}\Sigma_{jk}^{-1}+\hat{a}_{j}\hat{a}_{k}\Sigma_{il}^{-1}+\hat{a}_{j}\hat{a}_{l}\Sigma_{ik}^{-1}+\hat{a}_{k}\hat{a}_{l}\Sigma_{ij}^{-1}$
$+\Sigma_{ij}^{-1}\Sigma_{kl}^{-1}+\Sigma_{il}^{-1}\Sigma_{jk}^{-1}+\Sigma_{ik}^{-1}\Sigma_{jl}^{-1}$
@@
$\mathcal{C}=1+\frac{1}{24}\Phi_{n}^{-2}\sum_{IJKL}\kappa_{IJKL}\bigg{[}$ $\sum_{ijkl}\chi^{-1}_{iI}\chi^{-1}_{jJ}\chi^{-1}_{kK}\chi^{-1}_{lL}\bigg{(}\chi_{ij}\chi_{kl}+\chi_{il}\chi_{kj}+\chi_{ik}\chi_{jl}\bigg{)}$ (39)
$-\chi^{-1}_{IJ}\left(\sum_{kl}\chi_{kK}^{-1}\chi_{lL}^{-1}\chi_{kl}\right)-\chi^{-1}_{IK}\left(\sum_{jl}\chi_{jJ}^{-1}\chi_{lL}^{-1}\chi_{jl}\right)-\chi^{-1}_{IL}\left(\sum_{jk}\chi_{jJ}\chi_{kK}^{-1}\chi_{jk}\right)$
$-\chi^{-1}_{JK}\left(\sum_{il}\chi_{iI}^{-1}\chi_{lL}^{-1}\chi_{il}\right)-\chi^{-1}_{JL}\left(\sum_{ik}\chi_{iI}^{-1}\chi_{kK}^{-1}\chi_{ik}\right)-\chi^{-1}_{KL}\left(\sum_{ij}\chi_{iI}^{-1}\chi_{jJ}^{-1}\chi_{ij}\right)$
$+\chi^{-1}_{IJ}\chi^{-1}_{KL}+\chi^{-1}_{IK}\chi^{-1}_{JL}+\chi^{-1}_{IL}\chi^{-1}_{JK}\bigg{]}.$

### v2

@@ II.1 Gaussian noise
[…]
@@
[…]
@@ II.2 Non-Gaussian noise
$aligned p_{a}(a|\Phi,\vec{\alpha},\vec{\mu},\vec{c})&=\sum_{i=1}^{N}\alpha_{i}\frac{\exp\left\{-\frac{1}{2}\left[a-\mu_{i}\right]^{2}(c_{i}\Phi)^{-1}\right\}}{\sqrt{2\pi(c_{i}\Phi)}}\\ &=\sum_{i=1}^{N}\alpha_{i}\varphi(a,\mu_{i},\sqrt{c_{i}\Phi}),$ (4)
@@
$aligned\ln\int p(a,\Phi,\vec{\alpha},\vec{c}|\delta t)da=&\ln\sum_{i=1}^{N}\alpha_{i}\mathcal{L}(\delta t|c_{i}\Phi)\\ &+\ln\Pi(\Phi)+\ln\Pi(\vec{c})+\ln\Pi(\vec{\alpha}),$ (5)
@@ II.4 Non-Gaussian Likelihood
$aligned p_{n}(a|\Phi_{n},\alpha,c)&=(1-\alpha)\times\varphi(a,0,\sqrt{\Phi_{n}})+\alpha\times\varphi(a,0,\sqrt{c\Phi_{n}}),$ (13)
@@
$aligned\mathcal{L}(\delta t|\Phi^{c}_{n},\Phi^{s}_{n},\alpha,c,\vec{\lambda})&=(1-\alpha)^{2}\mathcal{L}(\delta t|\Phi_{n},\Phi_{n},\vec{\lambda})\\ &+\alpha^{2}\mathcal{L}(\delta t|c\Phi_{n},c\Phi_{n},\vec{\lambda})\\ &+(1-\alpha)\alpha\mathcal{L}(\delta t|\Phi_{n},c\Phi_{n},\vec{\lambda})\\ &+\alpha(1-\alpha)\mathcal{L}(\delta t|c\Phi_{n},\Phi_{n},\vec{\lambda}).$ (14)
@@ II.5 Free spectrum
When $a_{n}$ is described as a 2 Gaussian GMM, the result is given by Table 1. For more Gaussian in the GMM, we found no closed form expression of the ensemble average for the sum of the squares. However, the result can be easily obtained numerically by sampling the GMM. For a set of non-Gaussianity parameters $\Phi_{n}$, $\alpha$ and $c$, given that we have inferred the posterior distribution $p(\Phi_{n},\alpha,c|\delta t)$ from data $\delta t$, the probability distribution $p(a_{n},\Phi_{n},\vec{\alpha},\vec{c}|\delta t)$ is given by
@@
$aligned p(a_{n}|\delta t)&=\int p(a_{n},\Phi_{n},\alpha,c|\delta t)d\Phi_{n}d\alpha dc\\ &\approx\frac{1}{N_{D}}\sum_{j=1}^{N_{D}}p(a_{n}|\Phi_{j},\alpha_{j},c_{j}).$ (17)
@@
where the $a_{nj,i}^{c},a_{nj,i}^{s}$ are sampled from $p(a_{n},\Phi_{nj},\alpha_{j},c_{j}|\delta t)$ and $N_{s}$ is the total number of samples used to numerically estimate the average for the $j$-th posterior sample of $p(\Phi_{n},\alpha,c|\delta t)$. Here, we discard uneven moments because we only consider symmetric distributions. In this work, since we consider uncorrelated $a^{c}_{n}$ and $a^{s}_{n}$ modeled as a 2 Gaussian GMM, $\rho^{2}_{n,j}$ and $\rho^{4}_{n,j}$ can be directly obtained from the posterior samples of $\Phi_{nj},\alpha_{j},c_{j}$ using the expressions in Table 1.
@@ III Results
For each pulsar we consider a timing model accounting for the spin-down rate of that pulsar described by a quadratic polynomial in time and we marginalize over the first order errors in the timing model parameters [44, 2]. The full likelihood for common uncorrelated noise is written as a product of $N_{p}$ individual pulsar likelihoods given by Equation 14
$\mathcal{L}(\delta t|\Phi^{c}_{n},\Phi^{s}_{n},\alpha,c,\vec{\lambda})=\prod_{k=1}^{N_{p}}\mathcal{L}_{k}(\delta t_{k}|\Phi^{c}_{n},\Phi^{s}_{n},\alpha,c,\vec{\lambda}_{k})$ (21)
@@ III.1 Single pulsar noise
We test the model by injecting a time-correlated noise that has a consistent statistic across all frequency bins. In this example, we consider a noise that has Fourier coefficients $a$ distributed accroding to GMM of 2 Gaussians with $\alpha=0.5$ and $c=3$, following a powerlaw as in Equation 20 with $\log_{10}A=-13.5$ and $\gamma=3$ across 50 frequency bins that are the 100 first harmonics of $1/T_{obs}$. Recent works have shown the advantages of directly sampling the Fourier coefficients [29, 43]. The only disadvantange comes from the high number of coefficients there are to sample, but the recovered posterior for the unmarginalized parameters are identical to the traditionally marginalized likelihood.
The result of one simulation is shown in Figure 2 to verify that we indeed recover the injected values of the parameters. We only show the hyperparameters characterizing the noise, while numerically marginalizing over all Fourier coefficients by sampling them. We have used 50 frequencies, for a total of 100 Fourier coefficients that we fit for. This approach is straightforward as it requires no analytical marginalization. However, the sampling is affected by the high dimensionality of the parameter space and many samples are required to converge to the proper distribution. Still, in the case of GMMs, the non-marginalized (method I) and marginalized (method II) likelihoods presented in subsection II.4 should respond equally to non-Gaussianities present in the data. This can be seen in Figure 3 where we plot the Bayes factor $\mathcal{B}^{NG}_{G}$ comparing the non-Gaussian and Gaussian models as a function of the excess kurtosis $\Delta\bar{M}_{4}$. As $\Delta\bar{M}_{4}$ increases, the Bayes factor increases accordingly. Note that for $\Delta\bar{M}_{4}=0$ (i.e. Gaussian noise is injected), we have $\mathcal{B}^{NG}_{G}\approx 1$ meaning that the non-Gaussian model is not preferred over the Gaussian model.
@@ III.2 Common red noise
Figure 3: Bayes factor $\mathcal{B}^{NG}_{G}$ between the non-Gaussian and Gaussian models as a function of the injected relative excess kurtosis $\Delta\bar{M}_{4}$ calculated using Equation 7 for many simulations. The solid line is the median and the shaded region is bounded by the 16th and 84th percentiles. We show the two cases : (i) SPA (Single Pulsar Analysis) with 100 frequencies powerlaw red noise, (ii) CRN (Common red noise) with 100 pulsars.
@@ III.3 Background plus few sources
Figure 7: Excess moments as a function of the averaged SNR of the source for $N_{p}$ pulsars with $10$ years of observation, 100100 ns noise level and a GWB signal with $h_{c}=10^{-15}(f/f_{\rm yr})^{-2/3}$. Since $f_{\rm gw}=3/(10\textrm{yr})$, the source produces 3 cycles in the data. This plot shows the median excess moments estimated from a 100 realization of the CGW parameters. The shaded colored areas show the 16th and 84th percentiles of the realizations. Only the amplitude of the CGW was fixed to control the SNR.
@@ IV Conclusion
In this work, we introduced a method to model non-Gaussian features in PTA data by using GMMs as priors for the Fourier coefficients of stochastic processes. This approach enables analytical marginalization of the likelihood, making it computationally efficient and directly compatible with existing PTA analysis pipelines. We validated the method through simulations involving both single pulsar noise and common red noise processes, demonstrating that the model can robustly identify departures from Gaussianity using Bayesian model selection. The Bayes factor increases significantly in the presence of excess kurtosis, confirming the sensitivity of the approach to higher-order statistical structure. Furthermore, we applied the method to a more astrophysically motivated scenario involving a superposition of a Gaussian GWB and a few CGW sources. In this case, we showed that non-Gaussian features become detectable when a small number of loud sources dominate the signal, and that these features are suppressed once the deterministic CGWs are correctly modeled and subtracted. When a realistic GWB is generated from the superposition of the individual CGW emissions of a simulated population of SMBHBs, the non-Gaussianities are harder to detect and might require very large arrays of pulsars to increase their detectability. This is illustrated in Appendix  E where we show the Bayes factor $\mathcal{B}^{NG}_{G}$ for different sizes of array and varying excess kurtosis.
@@ Acknowledgements.
We thank Irene Ferranti for providing the realistic population catalogs. acknowledge support from the European Union’s H2020 ERC Advanced Grant “PINGU” (Grant No. 101142079) and from the PRIN 2022 grant ”GRAPE” (CUP H53D2300091 0006).
@@ Appendix A Moments of a Gaussian mixture
$aligned M_{n}&=\int(x-\alpha\mu_{0})^{n}p(x)dx\\ &=(1-\alpha)\int(x-\alpha\mu_{0})^{n}\varphi(x,0,\sqrt{\Phi})dx+\alpha\int(x-\alpha\mu_{0})^{n}\varphi(x,\mu_{0},\sqrt{c\Phi})dx\\ &=(1-\alpha)\int\tilde{x}^{n}\varphi(\tilde{x},-\alpha\mu_{0},\sqrt{\Phi})d\tilde{x}+\alpha\int\tilde{x}^{n}\varphi(\tilde{x},(1-\alpha)\mu_{0},\sqrt{c\Phi})d\tilde{x}\\ &=(1-\alpha)\mathcal{M}_{n}(-\alpha\mu_{0},\sqrt{\Phi})+\alpha\mathcal{M}_{n}((1-\alpha)\mu_{0},\sqrt{c\Phi}),$ (26)
@@ Appendix B Marginalized Gaussian likelihood
[…]
@@
$aligned\mathcal{L}(\delta t|a,\Phi,\mu)=&\frac{\exp\left\{-\frac{1}{2}\delta t^{\top}\left[N^{-1}-N^{-1}F\Sigma^{-1}F^{\top}N^{-1}\right]\delta t\right\}}{|2\pi N|^{1/2}|2\pi\Phi|^{1/2}}\times\exp\left\{-\frac{1}{2}\left[a-\hat{a}\right]^{\top}\Sigma\left[a-\hat{a}\right]\right\}\\ &\times\exp\left\{-\frac{1}{2}\mu^{\top}\left[\Phi^{-1}-\Phi^{-1}\Sigma^{-1}\Phi^{-1}\right]\mu+\mu^{\top}\Phi^{-1}\Sigma^{-1}F^{\top}N^{-1}\delta t\right\}$ (30)
@@
$aligned\mathcal{L}(\delta t|\Phi,\mu)&=\int_{-\infty}^{+\infty}da\mathcal{L}(\delta t|a,\Phi,\mu)\\ &=\frac{\exp\left\{-\frac{1}{2}\delta t^{\top}\left[N+F^{\top}\Phi F\right]^{-1}\delta t\right\}}{|2\pi N|^{1/2}|\Phi|^{1/2}|\Sigma|^{1/2}}\\ &\times\exp\left\{-\frac{1}{2}\mu^{\top}\left[\Phi^{-1}-\Phi^{-1}\Sigma^{-1}\Phi^{-1}\right]\mu+\mu^{\top}\Phi^{-1}\Sigma^{-1}F^{\top}N^{-1}\delta t\right\}$ (31)
@@ Appendix C Gram-Charlier A expansion for kurtosis
$aligned H_{IJKL}(\vec{a})=&(-1)^{4}\ \Pi(\vec{a})^{-1}\frac{\partial^{4}}{\partial a_{I}\partial a_{J}\partial a_{K}\partial a_{L}}\Pi(\vec{a})\\ =&\Phi^{-4}\left[(\sum_{i}\chi_{iI}^{-1}a_{i})(\sum_{j}\chi_{jJ}^{-1}a_{j})(\sum_{k}\chi_{kK}^{-1}a_{k})(\sum_{l}\chi_{lL}^{-1}a_{l})\right]\\ &-\Phi^{-3}\left[\chi^{-1}_{IJ}(\sum_{k}\chi_{kK}^{-1}a_{k})(\sum_{l}\chi_{lL}^{-1}a_{l})+\chi^{-1}_{IK}(\sum_{j}\chi_{jJ}^{-1}a_{j})(\sum_{l}\chi_{lL}^{-1}a_{l})+\chi^{-1}_{IL}(\sum_{j}\chi_{jJ}^{-1}a_{j})(\sum_{k}\chi_{kK}^{-1}a_{k})\right.\\ &-\left.\chi^{-1}_{JK}(\sum_{i}\chi_{iI}^{-1}a_{i})(\sum_{l}\chi_{lL}^{-1}a_{l})+\chi^{-1}_{JL}(\sum_{i}\chi_{iI}^{-1}a_{i})(\sum_{k}\chi_{kK}^{-1}a_{k})+\chi^{-1}_{KL}(\sum_{i}\chi_{iI}^{-1}a_{i})(\sum_{j}\chi_{jJ}^{-1}a_{j})\right]\\ &+\Phi^{-2}\bigg[\chi^{-1}_{IJ}\chi^{-1}_{KL}+\chi^{-1}_{IK}\chi^{-1}_{JL}+\chi^{-1}_{IL}\chi^{-1}_{JK}\bigg].$ (35)
@@
$aligned\mathcal{L}(\delta t|\Phi,\kappa)&=\int_{-\infty}^{+\infty}da\mathcal{L}(\delta t|a,\Phi,\kappa,\mu=0)\\ &=\mathcal{C}^{-1}\times\mathcal{L}(\delta t|\Phi,\kappa=0)\int da\frac{\exp\left\{-\frac{1}{2}\left[a-\hat{a}\right]^{\top}\Sigma\left[a-\hat{a}\right]\right\}}{|2\pi\Sigma^{-1}|}\left[1+\frac{1}{4!}\sum_{IJKL}\kappa_{IJKL}H_{IJKL}(\vec{a}_{n})\right]\\ &=\mathcal{C}^{-1}\times\mathcal{L}(\delta t|\Phi,\kappa=0)\left[1+\frac{1}{24}\sum_{IJKL}\kappa_{IJKL}\textrm{E}_{\Sigma}\left\{H_{IJKL}\right\}\right]$ (36)
@@
[…]
@@
[…]
@@
[…]